OpenStudy (anonymous):

Jason leaves Detroit at 3:00 PM and drives at a constant speed west along I-96. He passes Ann Arbor, 40 mi from Detroit, at 3:40 PM. (a) Express the distance traveled in terms of the time (in hours) elapsed.

OpenStudy (anonymous):

x=v*t 40=v*40/60 v=60